# DABL: Detecting Semantic Anomalies in Business Processes Using Large Language Models

### v1

@@ Abstract
Detecting anomalies in business processes is crucial for ensuring operational success. While many existing methods rely on statistical frequency to detect anomalies, it’s important to note that infrequent behavior doesn’t necessarily imply undesirability. To address this challenge, detecting anomalies from a semantic viewpoint proves to be a more effective approach. However, current semantic anomaly detection methods treat a trace (i.e., process instance) as multiple event pairs, disrupting long-distance dependencies. In this paper, we introduce DABL, a novel approach for detecting semantic anomalies in business processes using large language models (LLMs). We collect 143,137 real-world process models from various domains. By generating normal traces through the playout of these process models and simulating both ordering and exclusion anomalies, we fine-tune Llama 2 using the resulting log. Through extensive experiments, we demonstrate that DABL surpasses existing state-of-the-art semantic anomaly detection methods in terms of both generalization ability and learning of given processes. Users can directly apply DABL to detect semantic anomalies in their own datasets without the need for additional training. Furthermore, DABL offers the capability to interpret the causes of anomalies in natural language, providing valuable insights into the detected anomalies.
@@ Introduction
Business process anomaly detection is geared towards identifying undesired behavior occurring during process execution, serving as a crucial component in guaranteeing the efficient and dependable operation of businesses. By pinpointing anomalies within business processes, these detection techniques facilitate timely intervention, maintenance, and optimization, consequently bolstering overall well-being.
@@
Over the past few decades, notable advancements have been achieved in business process anomaly detection. Fig. 1 illustrates a comparison of various applicable methods for accomplishing this task. Traditional statistical-based approaches (Lu, Fang, and Fang 2022; Ko and Comuzzi 2022; Nolle et al. 2022; Guan et al. 2024) rely on analyzing statistical frequencies to identify anomalies. However, infrequent behavior is not necessarily anomalous; it may represent rare but acceptable behavior. Conversely, frequent behavior may not always be normal. Furthermore, these methods focus on providing anomaly scores and require manual specification of thresholds to distinguish between normal and anomalous instances, which is not suitable for real-world applications. Alternatively, methods based on conformance checking (Ebrahim and Golpayegani 2022; Sarno, Sinaga, and Sungkono 2020; Sinaga and Sarno 2016) detect anomalies by assessing the alignment between traces and their corresponding process models. Yet, accurately capturing complex processes within a process model remains a challenge, thereby restricting the utility of such approaches. The concept of semantic anomaly detection, recently introduced, addresses these challenges by identifying anomalies from a semantic viewpoint. For example, it can detect irregularities such as a claim being paid after having been rejected. Its grounding in natural language analysis allows for the consideration of typical behavior in standard processes, eliminating the necessity of having a specific process model at hand. However, existing semantic-based anomaly detection methods (van der Aa, Rebmann, and Leopold 2021; Caspary, Rebmann, and van der Aa 2023) treat a trace as multiple event pairs, disrupting long-distance dependencies and thus limiting their accuracy. Additionally, these methods interpret the cause of anomalies by providing anomalous event pairs, which can be confusing.
@@
Recently, there have been significant advancements in large language models (LLMs). Due to their remarkable language comprehension abilities, LLMs such as GPT-3.5 (Ouyang et al. 2022), GPT-4 (Achiam et al. 2023), Llama 2 (Touvron et al. 2023), and GLM-3 (Zeng et al. 2022) have shown proficiency in tasks like summarization, paraphrasing, and instruction following in zero-shot scenarios. However, in the context of semantic anomaly detection in business processes, their performance is limited by a lack of prior domain knowledge. As illustrated in Fig. 1, they often struggle to provide specific answers.
@@
To address the aforementioned issues, we propose DABL, a fine-tuned LLM designed to detect semantic anomalies in business processes. Due to the lack of event logs comprising traces from various domains with rich semantic information, we generated our training dataset by playing out 143,137 real-world process models from three different process model datasets. This resulted in 1,574,381 normal traces. The collected process models cover a broad range of domains, including common processes related to order and request handling, as well as specialized processes in fields such as software engineering and healthcare. Utilizing the generated normal traces, we then created synthetic anomalous traces. We introduced ordering anomalies, where activities should be executed in a different sequence (e.g., ”accept request” followed by ”check request”), and exclusion anomalies, where certain activities should not occur together within the same trace without an intermediate activity (e.g., ”refusing the application” followed by ”accepting the application” without ”reapplying” in between). These generated normal and anomalous traces collectively form the training dataset. Finally, by incorporating traces into question and answer content, we fine-tune the Llama 2-Chat 13B model (Touvron et al. 2023), an open-source LLM, using QLoRA (Dettmers et al. 2024), to create a generic model capable of detecting semantic anomalies in business processes. Compared to existing anomaly detection methods, DABL offers the capability to interpret the causes of anomalies in natural language, providing valuable insights into the detected anomalies. Extensive experiments show that DABL surpasses state-of-the-art methods in both generalization ability and learning of given processes. Thanks to its strong generalization ability, users can apply our open-source, fine-tuned model directly to their datasets without the need for additional training.
Our contributions are summarized as follows:
@@
  • We introduce DABL, an innovative method for fine-tuning large language models (LLMs) to detect semantic anomalies in business processes.
@@
  • Extensive experiments demonstrate that DABL outperforms state-of-the-art methods in both generalization ability and the learning of given processes. Tests on real-world datasets confirm the practical effectiveness of DABL.
@@
Some of these methods construct probabilistic models to infer the probability values (i.e., anomaly score) of traces. For example, HPDTMC (Yang et al. 2020) constructs discrete-time Markov chains (DTMC) and introduces hitting probabilities (HP). EDBN (Pauwels and Calders 2019a, b) extends dynamic Bayesian networks by adding functional dependencies. PN-BBN (Lu, Fang, and Fang 2022) extends Petri nets by incorporating Bayesian networks. Other methods convert traces into vector representations and then detect anomalies using data mining techniques such as local outlier factor (LOF) and isolation forest (IF). For example, activities are considered as words and encoded into vectors using the word2vec (Mikolov et al. 2013) technique in (Junior et al. 2020; Vertuam Neto et al. 2021). Additionally, Trace2vec (De Koninck, vanden Broucke, and De Weerdt 2018), extended from doc2vec and n-gram, is used to encode traces in (Rullo et al. 2020). Furthermore, the authors in (Ko and Comuzzi 2022, 2021) employ one-hot encoding to convert traces into vector representations and detect anomalies using statistical leverage (Hoaglin and Welsch 1978). Recently, deep learning has been adopted to detect anomalies based on the reconstruction error. Given that traces exhibit sequential data characteristics, the authors in (Guan et al. 2023; Nolle et al. 2022; Krajsic and Franczyk 2021) embed LSTM or GRU within the autoencoder to enhance the model’s reconstruction capabilities. The authors in (Huo et al. 2021; Guan et al. 2024; Niro and Werner 2024) transform traces into graphs and utilize graph neural networks (GNNs) to generate graph encodings, identifying anomalies by evaluating the reconstruction error of the graphs or traces.
@@
The conformance checking-based approaches (Ebrahim and Golpayegani 2022; Sarno, Sinaga, and Sungkono 2020; Sinaga and Sarno 2016) utilize process models, which are either provided by the user or derived from logs using process mining techniques. Anomalies are detected through conformance checking techniques (Leemans, Fahland, and van der Aalst 2018), which assess the alignment between traces and the corresponding process model. When the trace deviates from the process model, it is considered anomalous.
@@
Van der Aa et al. (van der Aa, Rebmann, and Leopold 2021) fine-tune BERT (Devlin et al. 2018), a pre-trained language model, to parse the names of executed activities by extracting the action and business object. Then, a knowledge base capturing assertions about the interrelations that should hold among actions parsed from names of executed activities is applied. The knowledge is extracted either from VerbOcean (Chklovski and Pantel 2004) or from an abstract representation of the process model. Anomalies can be detected by checking if the recorded process behavior violates the assertions captured in the knowledge base.
Caspary et al. (Caspary, Rebmann, and van der Aa 2023) extract event pairs that are in an eventually-follow relation. To detect anomalous event pairs, they propose two approaches: an SVM-based approach and a BERT-based approach. The SVM-based approach transforms an event pair into a vector representation using GloVe embeddings. This vector is then fed into a trained SVM, which classifies whether the event pair is an anomaly. The BERT-based approach extends BERT with an additional output layer for two-class classification, determining whether an input event pair is anomalous or not. Both the SVM and the extended BERT model are trained using normal event pairs extracted from normal traces, along with anomalous event pairs simulated by randomly generating event pairs that are not normal.
However, existing semantic-based methods treat a trace as multiple event pairs, which disrupts long-distance dependencies. Additionally, these methods only identify anomalous event pairs to interpret the causes of anomalies, making them difficult to understand. In contrast, our DABL incorporates the entire trace into a novel prompt, allowing the LLMs to automatically capture long-distance dependencies. DABL also provides insightful interpretations of the causes of anomalies in natural language, making them easy to understand.
@@
Motivated by the impressive cognitive abilities exhibited by large language models (LLMs) (Ouyang et al. 2022; Zeng et al. 2022; Achiam et al. 2023; Touvron et al. 2023), researchers have begun investigating their application for anomaly detection.
AnomalyGPT (Gu et al. 2024) and Myriad (Li et al. 2023) incorporate novel image encoders with LLMs for industrial anomaly detection (IAD). Elhafsi et al. (Elhafsi et al. 2023) apply an LLM to analyze potential confusion among observed objects in a scene, which could lead to task-relevant errors in policy implementation. LLMAD (Liu et al. 2024) leverages LLMs for few-shot anomaly detection by retrieving and utilizing both positive and negative similar time series segments. In (Qi et al. 2023) and (Egersdoerfer, Zhang, and Dai 2023), authors devise effective prompts to apply LLMs for zero/few-shot system log anomaly detection. SheepDog (Wu and Hooi 2023) conducts fake news detection by preprocessing data using LLMs to reframe the news, customizing each article to match different writing styles. Sarda et al. (Sarda et al. 2023) propose a pipeline for automatic microservice anomaly detection and remediation based on LLMs.
@@
DABL is a novel conversational fine-tuned large language model, primarily designed to detect semantic anomalies in business processes and interpret their causes. Fig. 2 details the DABL training procedure, which consists mainly of dataset preparation and fine-tuning. DABL is implemented in Python, and the source code is accessible at https://github.com/guanwei49/DABL.
@@
To effectively fine-tune LLMs for developing a generic model capable of detecting semantic anomalies in business processes, a log meeting the following criteria is imperative: i) it must encompass both normal and anomalous traces, ii) it should contain rich semantic information (i.e., the activities should not be represented by meaningless characters), and iii) the traces within it should stem from diverse processes across various domains. Since such a log is not available in the real world, we generate normal traces by playout of the real-world process models from the BPM Academic Initiative (BPMAI) (Weske et al. 2020), fundamentals of business process management (FBPM) (Dumas et al. 2018), and SAP signavio academic models (SAP-SAM) (Sola et al. 2022). These process models cover a broad range of domains, including common processes related to order and request handling, as well as specialized processes from fields such as software engineering and healthcare. We then generate synthetic anomalies from these normal traces. We detail the dataset preparation in the following subsection.
@@
We select process models $\mathcal{M}$ from BPMAI, FBPM, and SAP-SAM that meet the following criteria: they are in BPMN notation (Chinosi and Trombetta 2012), described in English, and convertible into a sound workflow net. This results in a total set of 144,137 process models. Among these, 143,137 process models are used for generating training datasets, resulting in 1,574,381 normal traces, while the remaining 1,000 process models are used for generating test datasets.
Next, for each process model $m\in\mathcal{M}$, we perform a playout to obtain the set of normal traces, denoted as $\mathcal{L}_{m}$. These traces are allowed by the process model $m$. To prevent infinite trace lengths, we limit each loop in the process model $m$ to be executed a maximum of twice.
@@
Ordering anomalies arise when activities ought to be executed in a different sequence. Five types of ordering anomalies, as identified in (Nolle et al. 2022), are frequently encountered in real-world business processes. Fig. 3 illustrates distinct anomalous traces resulting from the application of these five anomaly types to a normal trace $[e_{1},e_{2},\cdots,e_{n}]$. These anomaly types are defined as follows and are applied to a normal trace to generate ordering anomalies:
  • Skip: A sequence of up to three activities $[e_{i},\cdots,e_{j}]$ is skipped.
  • Insert: A sequence of up to three random activities $[e^{\prime}_{1},\cdots,e^{\prime}_{m}]$ is inserted. The random activities are selected from a set comprising all possible activities across all process models.
  • Rework: A sequence of up to three activities $[e_{i},\cdots,e_{j}]$ is executed a second time after activity $e_{k}$.
  • Early: A sequence of up to three activities $[e_{i},\cdots,e_{j}]$ is executed earlier and consequently skipped later.
  • Late: A sequence of up to three activities $[e_{i},\cdots,e_{j}]$ is executed later and consequently skipped earlier.
@@
  • Skip: The activities \${$e_{i},\cdots,e_{j}$} are skipped before \${$e_{j+1}$}.
  • Insert: The activities \${$e_{1}^{\prime},\cdots,e_{m}^{\prime}$} should not be executed.
  • Rework: The activities \${$e_{i},\cdots,e_{j}$} are reworked after \$$\$\{e_{k}\}$.
  • Early: The activities \${$e_{i},\cdots,e_{j}$} are executed too early, they should be executed after \${$e_{i-1}$}.
  • Late: The activities \${$e_{i},\cdots,e_{j}$} are executed too late, they should be executed before \${$e_{j+1}$}.
Here, \$$\{e_{i},\ldots,e_{j}\}$ represents converting the trace $[e_{i},\cdots,e_{j}]$ into a string format by enclosing each executed activity in apostrophes and separating them with commas, while using and before the penultimate and ultimate activities. For example, for the trace $[A,B,C,D]$, the resulting string would be ’A’, ’B’, ’C’ and ’D’.
@@
Figure 4: Different anomaly types applied to a normal trace.
The process tree (Aalst, Buijs, and Dongen 2011), a specialized form of process model, is utilized for analyzing process structure. For instance, the process tree corresponding to the process model depicted in Fig. 2 is illustrated in Fig. 4. We begin by converting the gathered process model $m$ into a process tree using the techniques presented in (van Zelst and Leemans 2020). Then, we replace an exclusive node (represented as $\times$) in the process tree with a parallel node (represented as $\wedge$), resulting in a modified process model denoted as $m^{\prime}$. This modification enables certain exclusive activities to be executed within the same trace to simulate exclusion anomalies. It is important to recognize that a single process tree may contain multiple exclusive nodes; therefore, we carry out this modification successively, resulting in multiple modified process models. Subsequently, we playout of all the modified models, restricting each loop to be executed a maximum of twice, to generate the set of traces, denoted as $\mathcal{L}_{m^{\prime}}$. We refine $\mathcal{L}_{m^{\prime}}$ by excluding traces present in $\mathcal{L}_{m}$ from it, resulting in the set of exclusion anomalies $\mathcal{L}_{m}^{e}$.
[…]
Formally, consider a modified model $m^{\prime}$, which results from modifying an exclusive node $R$ in the process tree corresponding to the process model $m$. An exclusion anomaly $t$ is generated from model $m^{\prime}$. The node $R$ has $N$ branches, with the activity sets $\mathcal{A}_{1},\cdots,\mathcal{A}_{N}$ under them. Activities within each activity set $\mathcal{A}_{i}$ that do not appear in $t$ are filtered out. The cause of this exclusion anomaly $t$ is then:
  • The activities \$$\mathcal{A}_{1}$} are mutually exclusive with the activities \$$\bigcup_{i=2\cdots N}\mathcal{A}_{i}$}, meaning they should not be executed within the same process instance.
@@
The first component introduces the traces, such as ”In the following business process trace, each executed activity is separated by a comma: [Send acceptance pack, Check credit history, Assess loan risk, Assess eligibility, Prepare acceptance pack]”. The second component queries whether the trace is anomalous, asking, for instance, ”Is this trace normal or anomalous?”. The LLM first responds to whether the given trace is normal or anomalous. If it is anomalous, the LLM is asked about the cause of the anomaly, for example, ”What causes this trace to deviate?”. The LLM then explains the cause of the anomaly, such as ”The activity ’Send acceptance pack’ is executed too early, it should be executed after ’Prepare acceptance pack’.”. This descriptive content about the anomaly’s cause provides valuable insights and facilitates actions to maintain the health of the process execution.
@@
DABL involves fine-tuning the open-source Llama 2-Chat 13B model (Touvron et al. 2023) to enhance its capability to detect semantic anomalies in business processes. To mitigate the expense associated with fine-tuning LLMs with a substantial parameter count, we leverage QLoRA (Dettmers et al. 2024) to reduce memory usage. QLoRA achieves this by back-propagating gradients into a frozen 4-bit quantized model while preserving the performance level attained during the full 16-bit fine-tuning process.
We employ the Adam optimizer (Kingma and Ba 2014) to fine-tune the LLMs for two epochs, setting the initial learning rate to $5\times 10^{-5}$ with polynomial learning rate decay. The mini-batch size is set to 64. The fine-tuning is carried out on an NVIDIA A6000 GPU with 48 GB of memory.
@@
As mentioned in previous section, we allocate 1,000 process models for generating the test dataset $\mathcal{D}_{1}$. These models produce 14,387 normal traces, and we randomly simulate anomalies, resulting in 13,694 anomalous traces. In total, the test dataset $\mathcal{D}_{1}$ comprises 28,081 traces.
@@
Compared Methods
In our evaluation, we compare our DABL to existing semantic business process anomaly detection methods: SENSE (Caspary, Rebmann, and van der Aa 2023) and SEM (van der Aa, Rebmann, and Leopold 2021). SENSE offers both SVM-based and BERT-based models for detecting anomalous event pairs, which we denote as SENSE-SVM and SENSE-BERT, respectively. These methods divide traces into event pairs and determine whether each pair is normal or anomalous. If at least one event pair in a trace is identified as anomalous, the entire trace is classified as anomalous. It is important to note that SEM can only detect anomalous event pairs that share the same business object, automatically classifying pairs with distinct business objects as non-anomalous. Due to the high training costs, we utilize the open-source trained models provided by the authors for the test dataset $\mathcal{D}_{1}$. For the test dataset $\mathcal{D}_{2}$, we train the comparative models using the 21,298 normal traces available within it. The hyper-parameters of these methods are set to the values that yielded the best results reported in the original paper.
Evaluation Metrics
Following existing anomaly detection methods, we employ precision, recall, F1-score and accuracy to evaluate the anomaly detection performance.
The recall-oriented understudy for gisting evaluation (ROUGE) (Lin 2004) is a software package and metric set designed to assess the quality of generated text by comparing it with ground truth text. In our evaluation of DABL’s ability to interpret the cause of anomalies, we utilize ROUGE-2 and ROUGE-L metrics.
Quantitative Results
Anomaly Detection

### v2

@@ Abstract
Detecting anomalies in business processes is crucial for ensuring operational success. While many existing methods rely on statistical frequency to detect anomalies, it’s important to note that infrequent behavior doesn’t necessarily imply undesirability. To address this challenge, detecting anomalies from a semantic viewpoint proves to be a more effective approach. However, current semantic anomaly detection methods treat a trace (i.e., process instance) as multiple event pairs, disrupting long-distance dependencies. In this paper, we introduce DABL, a novel approach for detecting semantic anomalies in business processes using large language models (LLMs). We collect 143,137 real-world process models from various domains. By generating normal traces through the playout of these process models and simulating both ordering and exclusion anomalies, we fine-tune Llama 2 using the resulting log. Through extensive experiments, we demonstrate that DABL surpasses existing state-of-the-art semantic anomaly detection methods in terms of both generalization ability and learning of given processes. Users can directly apply DABL to detect semantic anomalies in their own datasets without the need for additional training. Furthermore, DABL offers the ability to interpret anomalies’ causes in natural language, providing valuable insights into the detected anomalies.
Codehttps://github.com/guanwei49/DABL
@@ Introduction
Business process anomaly detection is geared towards identifying undesired behavior occurring during process execution, serving as a crucial component in guaranteeing the efficient and dependable operation of businesses. By pinpointing anomalies within business processes, these detection techniques facilitate timely intervention, maintenance, and optimization, consequently bolstering overall well-being (10738505).
@@
Over the past few decades, notable advancements have been achieved in business process anomaly detection. Fig. 1 illustrates a comparison of various applicable methods for accomplishing this task. Traditional statistical-based approaches (Lu2022; Ko2022; nolle2022binet; Guan2023) rely on analyzing statistical frequencies to identify anomalies. However, infrequent behavior is not necessarily anomalous; it may represent rare but acceptable behavior. Conversely, frequent behavior may not always be normal. Furthermore, these methods focus on providing anomaly scores and require manual specification of thresholds to distinguish between normal and anomalous instances, which is not suitable for real-world applications. Alternatively, methods based on conformance checking (DBLP:journals/virology/EbrahimG22; DBLP:journals/jbd/SarnoSS20; sinaga2016business) detect anomalies by assessing the alignment between traces and their corresponding process models. Yet, accurately capturing complex processes within a process model remains a challenge, thereby restricting the utility of such approaches. The concept of semantic anomaly detection, recently introduced, addresses these challenges by identifying anomalies from a semantic viewpoint. For example, it can detect irregularities such as a claim being paid after having been rejected. Its grounding in natural language analysis allows for the consideration of typical behavior in standard processes, eliminating the necessity of having a specific process model at hand. However, existing semantic-based anomaly detection methods (van2021natural; caspary2023does) treat a trace as multiple event pairs, disrupting long-distance dependencies and thus limiting their accuracy. Additionally, these methods interpret the cause of anomalies by providing anomalous event pairs, which can be confusing.
@@
Recently, there have been significant advancements in LLMs. Due to their remarkable language comprehension abilities, LLMs such as GPT-3.5 (ouyang2022training), GPT-4 (achiam2023gpt), Llama 2 (touvron2023llama), and GLM-3 (zeng2022glm) have shown proficiency in tasks like summarization, paraphrasing, and instruction following in zero-shot scenarios. However, in the context of semantic anomaly detection in business processes, their performance is limited by a lack of prior domain knowledge. As illustrated in Fig. 1, they often struggle to provide specific answers.
@@
To address the aforementioned issues, we propose DABL, a fine-tuned LLM designed to detect semantic anomalies in business processes. Due to the lack of event logs comprising traces from various domains with rich semantic information, we generated our training dataset by playing out 143,137 real-world process models from three different process model datasets. This resulted in 1,574,381 normal traces. The collected process models cover a broad range of domains, including common processes related to order and request handling, as well as specialized processes in fields such as software engineering and healthcare. Utilizing the generated normal traces, we then created synthetic anomalous traces. We introduced ordering anomalies, where activities should be executed in a different sequence (e.g., ”accept request” followed by ”check request”), and exclusion anomalies, where certain activities should not occur together within the same trace without an intermediate activity (e.g., ”refusing the application” followed by ”accepting the application” without ”reapplying” in between). These generated normal and anomalous traces collectively form the training dataset. Finally, by incorporating traces into question and answer content, we fine-tune the Llama 2-Chat 13B model (touvron2023llama), an open-source LLM, using QLoRA (dettmers2024qlora), to create a generic model capable of detecting semantic anomalies in business processes. Compared to existing anomaly detection methods, DABL offers the capability to interpret the causes of anomalies in natural language, providing valuable insights into the detected anomalies. Extensive experiments show that DABL surpasses state-of-the-art methods in both generalization ability and learning of given processes.
Thanks to DABL’s strong generalization ability, users can apply our open-source, fine-tuned model directly to their datasets without the need for additional training. Notably, it operates in a zero-shot manner, meaning it does not need normal traces or a process model during its operation. Our contributions are summarized as follows:
@@
  • We introduce DABL, an innovative method for fine-tuning LLMs to detect semantic anomalies in business processes.
@@
  • Extensive experiments show that DABL outperforms state-of-the-art methods in both generalization ability and the learning of given processes. Tests on real-world datasets confirm the practical effectiveness of DABL.
@@
Some of these methods construct probabilistic models to infer the probability values (i.e., anomaly score) of traces. For example, HPDTMC (9408048) constructs discrete-time Markov chains (DTMC) and introduces hitting probabilities (HP). EDBN (10.1145/3297280.3297326; 10.1145/3357385.3357387) extends dynamic Bayesian networks by adding functional dependencies. PN-BBN (Lu2022) extends Petri nets by incorporating Bayesian networks. Other methods convert traces into vector representations and then detect anomalies using data mining techniques such as local outlier factor (LOF) and isolation forest (IF). For example, activities are considered as words and encoded into vectors using the word2vec (DBLP:journals/corr/abs-1301-3781) technique in (9230308; 10.1145/3466933.3466979). Additionally, Trace2vec (10.1007/978-3-319-98648-7_18), extended from doc2vec and n-gram, is used to encode traces in (rullo2020framework). Furthermore, the authors in (Ko2022; Ko202153) employ one-hot encoding to convert traces into vector representations and detect anomalies using statistical leverage (hoaglin1978hat). Recently, deep learning has been adopted to detect anomalies based on reconstruction errors. Given that traces exhibit sequential data characteristics, the authors in (10088425; nolle2022binet; DBLP:conf/kes/KrajsicF21; guan2024comb) embed LSTM, GRU or transformers within autoencoders to enhance the model’s reconstruction capabilities. The authors in (DBLP:conf/bpm/HuoVRAIM21; Guan2023; niro2024detecting) transform traces into graphs and utilize graph neural networks (GNNs) to generate graph encodings, identifying anomalies by evaluating the reconstruction error of the graphs or traces.
@@
The conformance checking-based approaches (DBLP:journals/virology/EbrahimG22; DBLP:journals/jbd/SarnoSS20; sinaga2016business) utilize process models, which are either provided by the user or derived from logs using process mining techniques. Anomalies are detected through conformance checking techniques (DBLP:journals/sosym/LeemansFA18), which assess the alignment between traces and the corresponding process model. When the trace deviates from the process model, it is considered anomalous.
@@
Van der Aa et al. (van2021natural) fine-tune BERT (devlin2018bert), a pre-trained language model, to parse the names of executed activities by extracting the action and business object. Then, a knowledge base capturing assertions about the interrelations that should hold among actions parsed from names of executed activities is applied. The knowledge is extracted either from VerbOcean (chklovski2004verbocean) or from an abstract representation of the process model. Anomalies can be detected by checking if the recorded process behavior violates the assertions captured in the knowledge base.
Caspary et al. (caspary2023does) extract event pairs that are in an eventually-follow relation. To detect anomalous event pairs, they propose two approaches: an SVM-based approach and a BERT-based approach. The SVM-based approach transforms an event pair into a vector representation using GloVe embeddings. This vector is then fed into a trained SVM, which classifies whether the event pair is an anomaly. The BERT-based approach extends BERT with an additional output layer for two-class classification, determining whether an input event pair is anomalous or not. Both the SVM and the extended BERT model are trained using normal event pairs extracted from normal traces, along with anomalous event pairs simulated by randomly generating event pairs that are not normal.
However, existing semantic-based methods treat a trace as multiple event pairs, which disrupts long-distance dependencies. Additionally, these methods only identify anomalous event pairs to interpret the causes of anomalies, making them difficult to understand. In contrast, our DABL incorporates the entire trace into a novel prompt, allowing the LLMs to capture long-distance dependencies. DABL also provides insightful interpretations of the causes of anomalies in natural language, making them easy to understand.
@@
Motivated by the impressive cognitive abilities exhibited by LLMs (ouyang2022training; zeng2022glm; achiam2023gpt; touvron2023llama), researchers have begun investigating their application for anomaly detection. AnomalyGPT (gu2024anomalygpt) and Myriad (li2023myriad) incorporate novel image encoders with LLMs for industrial anomaly detection (IAD). Elhafsi et al. (elhafsi2023semantic) apply an LLM to analyze potential confusion among observed objects in a scene, which could lead to task-relevant errors in policy implementation. LLMAD (liu2024large) leverages LLMs for few-shot anomaly detection by retrieving and utilizing both positive and negative similar time series segments. In (qi2023loggpt; guan2024logllm; egersdoerfer2023early), authors explore the application of LLMs for log-based anomaly detection. SheepDog (wu2023fake) conducts fake news detection by preprocessing data using LLMs to reframe the news, customizing each article to match different writing styles. Sarda et al. (sarda2023adarma) propose a pipeline for automatic microservice anomaly detection and remediation based on LLMs.
@@
DABL is a novel conversational fine-tuned large language model, primarily designed to detect semantic anomalies in business processes and interpret their causes. Fig. 2 details the DABL training procedure, which consists mainly of dataset preparation and fine-tuning.
@@
To effectively fine-tune LLMs for developing a generic model capable of detecting semantic anomalies in business processes, a log meeting the following criteria is imperative: i) it must encompass both normal and anomalous traces, ii) it should contain rich semantic information (i.e., the activities should not be represented by meaningless characters), and iii) the traces within it should stem from diverse processes across various domains. Since such a log is not available in the real world, we generate normal traces by playout of the real-world process models from the BPM Academic Initiative (BPMAI) (Weske2020ModelCO), fundamentals of business process management (FBPM) (dumas2018fundamentals), and SAP signavio academic models (SAP-SAM) (sola2022sap). These process models cover a broad range of domains, including common processes related to order and request handling, as well as specialized processes from fields such as software engineering and healthcare. We then generate synthetic anomalies from these normal traces. We detail the dataset preparation in the following subsection.
@@
We select process models $\mathcal{M}$ from BPMAI, FBPM, and SAP-SAM that meet the following criteria: they are in BPMN notation (chinosi2012bpmn), described in English, and convertible into a sound workflow net (van2011soundness). This results in a total set of 144,137 process models. Among these, 143,137 process models are used for generating training datasets, resulting in 1,574,381 normal traces, while the remaining 1,000 process models are used for generating test datasets.
Next, for each process model $m\in\mathcal{M}$, we perform a playout (berti2019process) to obtain the set of normal traces, denoted as $\mathcal{L}_{m}$. These traces are allowed by the process model $m$. To prevent infinite trace lengths, we limit each loop in the process model $m$ to be executed a maximum of twice.
@@
Ordering anomalies arise when activities ought to be executed in a different sequence. Five types of ordering anomalies, as identified in (nolle2022binet), are frequently encountered in real-world business processes. Fig. 3 illustrates distinct anomalous traces resulting from the application of these five anomaly types to a normal trace $\langle e_{1},e_{2},\cdots,e_{n}\rangle$. These anomaly types are defined as follows and are applied to a normal trace to generate ordering anomalies:
  • Skip: A sequence of up to three activities $\langle e_{i},\cdots,e_{j}\rangle$ is skipped.
  • Insert: A sequence of up to three random activities $\langle e^{\prime}_{1},\cdots,e^{\prime}_{m}\rangle$ is inserted. The random activities are selected from a set comprising all possible activities across all process models.
  • Rework: A sequence of up to three activities $\langle e_{i},\cdots,e_{j}\rangle$ is executed a second time after activity $e_{k}$.
  • Early: A sequence of up to three activities $\langle e_{i},\cdots,e_{j}\rangle$ is executed earlier and consequently skipped later.
  • Late: A sequence of up to three activities $\langle e_{i},\cdots,e_{j}\rangle$ is executed later and consequently skipped earlier.
@@
  • Skip: The activities $e_{i},\cdots,e_{j}$} are skipped before $e_{j+1}$}.
  • Insert: The activities $e_{1}^{\prime},\cdots,e_{m}^{\prime}$} should not be executed.
  • Rework: The activities $e_{i},\cdots,e_{j}$} are reworked after $\mathdollar\{e_{k}\}$.
  • Early: The activities $e_{i},\cdots,e_{j}$} are executed too early, they should be executed after $e_{i-1}$}.
  • Late: The activities $e_{i},\cdots,e_{j}$} are executed too late, they should be executed before $e_{j+1}$}.
[…]
@@
Figure 4: The process tree corresponding to the process model depicted in Fig. 2.
The process tree (aalst2011towards), a specialized form of process model, is utilized for analyzing process structure. For instance, the process tree corresponding to the process model depicted in Fig. 2 is illustrated in Fig. 4. We begin by converting the gathered process model $m$ into a process tree using the techniques presented in (van2020translating). Then, we replace an exclusive node (represented as $\times$) in the process tree with a parallel node (represented as $\wedge$), resulting in a modified process model denoted as $m^{\prime}$. This modification enables certain exclusive activities to be executed within the same trace to simulate exclusion anomalies. It is important to recognize that a single process tree may contain multiple exclusive nodes; therefore, we carry out this modification successively, resulting in multiple modified process models. Subsequently, we playout of all the modified models, restricting each loop to be executed a maximum of twice, to generate the set of traces, denoted as $\mathcal{L}_{m^{\prime}}$. We refine $\mathcal{L}_{m^{\prime}}$ by excluding traces present in $\mathcal{L}_{m}$ from it, resulting in the set of exclusion anomalies $\mathcal{L}_{m}^{e}$.
To extract the causes of exclusion anomalies, we need to identify activities that exhibit exclusion relationships. In a process tree, activities located under different branches of an exclusive node (represented as $\times$) exhibit such relationships. For example, in the process tree illustrated in Fig. 4, the activity set {Prepare acceptance pack, Send acceptance pack} and the activity set {Reject application} exhibit exclusion relationships. Formally, consider a modified model $m^{\prime}$, which results from modifying an exclusive node $R$ in the process tree corresponding to the process model $m$. An exclusion anomaly $t$ is generated from model $m^{\prime}$. The node $R$ has $N$ branches, with the activity sets $\mathcal{A}_{1},\cdots,\mathcal{A}_{N}$ under them. Activities within each activity set $\mathcal{A}_{i}$ that do not appear in $t$ are filtered out. The cause of this exclusion anomaly $t$ is then:
  • The activities $\mathcal{A}_{1}$} are mutually exclusive with the activities $\bigcup_{i=2\cdots N}\mathcal{A}_{i}$}, meaning they should not be executed within the same process instance.
@@
[…]
@@
DABL involves fine-tuning the open-source Llama 2-Chat 13B model (touvron2023llama) to enhance its capability to detect semantic anomalies in business processes. To mitigate the expense associated with fine-tuning LLMs with a substantial parameter count, we leverage QLoRA (dettmers2024qlora) to reduce memory usage. QLoRA achieves this by back-propagating gradients into a frozen 4-bit quantized model while preserving the performance level attained during the full 16-bit fine-tuning process.
We employ the Adam optimizer (kingma2014adam) to fine-tune the LLMs for two epochs, setting the initial learning rate to $5\times 10^{-5}$ with polynomial learning rate decay. The mini-batch size is set to 64. The fine-tuning is carried out on an NVIDIA A6000 GPU with 48 GB of memory.
@@
As mentioned in the previous section, we allocate 1,000 process models for generating the test dataset $\mathcal{D}_{1}$. These models produce 14,387 normal traces, and we randomly simulate anomalies, resulting in 13,694 anomalous traces. In total, the test dataset $\mathcal{D}_{1}$ comprises 28,081 traces.